# Ricci almost solitons: complete examples

### v1

@@ 1. Introduction and results
${\rm Ric}+\frac{1}{2}\mathcal{L}_{X}g+\lambda g=0$
for the Lie derivative in the direction of some vector field $X$ and a constant $\lambda$. Any Einstein space is trivially a Ricci soliton with $X=0$. If in addition $X$ is assumed to be a non-trivial conformal vector field, then it immediately follows that the Ricci soliton is Einstein. This case is often discussed in the literature, see [11], [6]. Besides the standard sphere, compact Ricci solitons seem to be rare [4, 5]. Sophisticated analytic and topological methods have been employed for obtaining non-compact Ricci solitons [22], [3], for the approach by doubly warped products see [10].
@@
for a constant $\rho\in\mathbb{R}$ is said to evolve by the Ricci-Bourguignon flow. Here we will consider two particular cases. For $\rho=0$ this is the Ricci flow, for $\rho=1/n$ the right hand side equals the traceless Ricci tensor. This flow was introduced by Bourguignon [2, 3.24].
@@
Ricci-Bourguignon gradient solitons are also named gradient $\rho$-Einstein solitons, if $\lambda$ in Equation (2) is allowed to be a function, then the metric is also called gradient Ricci-Bourguignon almost soliton. For $\lambda=0$ the gradient Ricci-Bourguignon (almost) soliton is called steady.
@@
All complete warped products with a $1$-dimensional basis satifying the equation $\textrm{Ric}+\nabla^{2}f-\frac{S}{n}g=0$ of a steady Ricci-Bourguignon soliton can be classified by a family of explicitly given functions $h$ and $f$ depending on finitely many parameters.
@@ 2. Basic formulas
and a function $f(t)$ defined on the basis. Unless stated otherwise, it will be assumed that $h$ and $f$ are smooth functions and $h^{\prime}$ is not constant. Then the gradients of $h$ and $f$ appear as $h^{\prime}\partial_{t}$ and $f^{\prime}\partial_{t}$, respectively, and the Hessian tensors are $\nabla^{2}h=h^{\prime\prime}g$ on the one hand and
@@
The Ricci tensors of $g$ and $g_{*}$ will be denoted by Ric and ${\rm Ric}_{*}$. Any equation of the type
@@
$a\textrm{Ric}(\partial_{t},\partial_{t})+bf^{\prime\prime}+c$ $=$ 00 0
@@
(4) $-(n-1)a\frac{h^{\prime\prime\prime}}{h^{\prime}}+bf^{\prime\prime}+c=0$
@@
Therefore one can express $f^{\prime}$ as $h^{\prime}$ times an integral of $\frac{a}{b}(n-2)\frac{h^{\prime\prime 2}-h^{\prime}h^{\prime\prime\prime}-\lambda_{*}}{h^{\prime 3}}$ and determine $f$ itself by an additional integration step, compare [18, Ex.1.5]. Finally, $c$ is obtained by the trace of $a{\rm Ric}+b\nabla^{2}f$. However, in general $c$ won’t be constant, so this method does not produce Ricci solitons but only Ricci almost solitons. Considering only the traceless part leads to the following:
@@
For any given Einstein space $M_{*}$ and any smooth function $h^{\prime}$ there is a function $f$ such that the warped product ×$\mathbb{R}\times_{h^{\prime}}M_{*}$ admits at least locally a solution of $(a{\rm Ric}+b\nabla^{2}f)^{\circ}=0$. For $a=b=1$ the left hand side is nothing but the traceless Bakry-Émery-Ricci tensor [21].
If $h^{\prime}$ is globally defined and everywhere strictly positive, then the warped product ×$\mathbb{R}\times_{h^{\prime}}M_{*}$ is complete whenever $M_{*}$ is complete. This leads to complete solutions of $({\rm Ric}+\nabla^{2}f)^{\circ}=0$ (Ricci almost solitons).
@@
    Essentially the same holds under the assumption that $a=b=1$ and $M$ is Einstein, compare [18, Thm.1.3]. Up to scaling, the list of Einstein spaces admitting a function $f$ satisfying $\nabla^{2}f=\lambda g$ for some function $\lambda$ consists of warped products $dt^{2}+(f^{\prime}(t))^{2}ds_{*}^{2}$ with a $1$-dimensional basis where $f(t)$ is a solution of the oscillator equation $f^{\prime\prime}+\sigma f=0$ where $\sigma$ is a normalized scalar curvature ($\sigma=1$ for the unit sphere) [13], [14]. Among them the unit sphere is the only compact Einstein space with $f^{\prime}(t)=\sin t$ with the two zeros added (it is not a global warped product).
@@
A solution with $a=b=1$ and a constant $c$ is normally called a gradient Ricci soliton, and the tensor ${\rm Ric}^{\infty}_{f}={\rm Ric}+\nabla^{2}f$ is called the Bakry-Émery-Ricci Tensor [21], for ${\rm Ric}^{m}_{f}$ see [9]. A solution with $a=b=1$ and a function $c$ is called a gradient Ricci almost soliton [18].
@@
    Any $2$-dimensional gradient Ricci soliton with $b\not=0$ is a warped product outside the set of critical points which are isolated. This follows from the known solutions of the equation $\nabla^{2}f=\phi g$ with any scalar function $\phi$ [19], [13], compare [17].
@@
1. The metric $ds^{2}=dt^{2}+\tanh^{2}t\,du^{2}$ on $(t,u)\in[0,\infty)\times S^{1}$ is a well known example of a $2$-dimensional complete Ricci soliton defining a complete metric on $\mathbb{R}^{2},$ cf. Kühnel [13, Example 4(a)]. Here $c=0,\gamma=-2,h^{\prime}(t)=\tanh t,f^{\prime}=-2h^{\prime},h^{\prime\prime}(t)=(\cosh t)^{-2},K=-h^{\prime\prime\prime}/h^{\prime}=2h^{\prime\prime}=-f^{\prime\prime},\nabla^{2}f=f^{\prime\prime}g=-2h^{\prime\prime}g=K\,g.$ With $x=\sinh t\,\sin u;y=\sinh t\cos u$ this metric on $\mathbb{R}^{2}$ is of the form
@@ 3. Complete and explicit solutions
If $n=2$ then the two coupled equations (4) und (6) read as:
$-ah^{\prime\prime\prime}+bh^{\prime}f^{\prime\prime}+ch^{\prime}$ $=$ 00 0
$ah^{\prime}h^{\prime\prime\prime}-bf^{\prime}h^{\prime}h^{\prime\prime}-ch^{\prime 2}$ $=$ 00 0
@@
(11) $-ah^{\prime\prime\prime}+b\gamma h^{\prime}h^{\prime\prime}+ch^{\prime}$ $=$ 00 0
@@
Proof. Both equations are simultaneously satisfied if and only if either $b=0$ or $h^{\prime}f^{\prime\prime}=h^{\prime\prime}f^{\prime}$ which in turn is equivalent to $\log(f^{\prime})^{\prime}=\log(h^{\prime})^{\prime}$ or $f^{\prime}=\gamma h^{\prime}$ for a constant $\gamma$. $\Box$
There is a method for solving an ODE of type $y^{\prime\prime}=F(y,y^{\prime})$ [20, §11]. We concentrate on the case $a=b=1$.
@@
For $c=0$ Equation (11) is of first order by $(h^{\prime\prime}-\frac{\gamma}{2}h^{\prime 2})^{\prime}=0$. There are various odd functions $h^{\prime}$ solving this equation. Typical solutions with initial condition $h^{\prime}(0)=0$ are $h^{\prime}(t)=\tan t$ for $\gamma=2$ and $h^{\prime}(t)=\tanh t=\frac{e^{2t}-1}{e^{2t}+1}$ for $\gamma=-2$ (the complete cigar metric. Other typical solutions are $h^{\prime}(t)=\frac{1}{t}$ for $\gamma=-2$ and $h^{\prime}(t)=\coth t=\frac{e^{2t}+1}{e^{2t}-1}$ for $\gamma=-2$. There are no even functions solving this equation.
@@
Let $h(t)=\arctan(t)$, then for any $\lambda_{*}$ there is a global function $f(t)$ such that ×$\mathbb{R}\times_{h^{\prime}}M_{*}$ is a complete Ricci almost soliton whenever $M_{*}$ is complete with Einstein constant $\lambda_{*}$.
@@
[…]
$=\frac{a\lambda_{*}}{b}\Big{(}\frac{t^{7}}{7}+\frac{3t^{5}}{5}+t^{3}+t\Big{)}+\frac{a(n-2)}{b}\Big{(}2t-4\arctan(t)+C\Big{)}.$
@@
[…]
@@
Let $h(t)=\frac{1}{3}t^{3}+t$, then for any $\lambda_{*}$ there is a global function $f(t)$ such that ×$\mathbb{R}\times_{h^{\prime}}M_{*}$ is a complete Ricci almost soliton whenever $M_{*}$ is complete with Einstein constant $\lambda_{*}$.
@@
Let $h(t)=(t^{2}+2)e^{t}$, then for any $\lambda_{*}$ there is a global function $f(t)$ such that ×$\mathbb{R}\times_{h^{\prime}}M_{*}$ is a complete Ricci almost soliton whenever $M_{*}$ is complete with Einstein constant $\lambda_{*}$.
@@
$=-2a(n-2)(t^{2}+12t+8)e^{2t}+b(t^{2}+2t+2)^{3}e^{3t}\Big{(}\frac{f^{\prime}}{(t^{2}+2t+2)e^{t}}\Big{)}^{\prime}=a\lambda_{*}.$
@@ 4. Ricci-Bourguignon solitons
with a constant $\lambda$ where $S$ denotes the scalar curvature [6]. The case $\lambda=0$ is then referred to as a steady Ricci-Bourguignon gradient soliton.
@@
2. If the gradient of $f$ is assumed to be a conformal vector field, then Equation (12) implies that $M$ is Einstein. Consequently, even if $\lambda$ is admitted to be a non-constant function, the only complete cases are listed in [13, Thm.27], implicitly already in [19, Sect.3]. In particular, the only compact case is the standard sphere. Implicitly this result can already be found in [19, Thm.1] since an Einstein space conformal to the standard sphere is isometric to the standard sphere. It is restated in [6, Thm.1.5]. For the last statement a constant scalar curvature is sufficient [13, Thm.24].
3. Complete gradient Ricci-Bourguignon almost solitons can be constructed by the same principle as in Section 3, compare [6, Ex.1.7].
@@
A warped product $M=\mathbb{R}\times_{h^{\prime}}M_{*}$ of dimension $n\geq 4$ with a potential function $f(t)$ is a steady Ricci-Bourguignon gradient soliton if and only if $f$ is harmonic and Equation (8) is satisfied with $a=b=1$.
@@
$ds^{2}=dt^{2}+(h^{\prime}(t))^{2}ds_{*}^{2}$ is a steady Ricci-Bourguignon gradient soliton if and only if $h$ and $f$ satisfy the following equations with the constant $\lambda_{*}$ as the normalized scalar curvature of $g_{*}$:
@@
(14) $\Delta f=f^{\prime\prime}+(n-1)\frac{f^{\prime}h^{\prime\prime}}{h^{\prime}}$ $=$ $0$
@@
$(n-2)\big{[}h^{\prime\prime 2}-h^{\prime}h^{\prime\prime\prime}\big{]}+h^{\prime}\big{[}f^{\prime\prime}h^{\prime}-f^{\prime}h^{\prime\prime}\big{]}$ $=$ 00 0
$f^{\prime\prime}+(n-1)\frac{f^{\prime}h^{\prime\prime}}{h^{\prime}}$ $=$ $0$

### v2

@@ 1. Introduction and results
${\rm Ric}+\frac{1}{2}\mathcal{L}_{X}g=\lambda g$
for the Lie derivative $\mathcal{L}_{X}g$ in the direction of some vector field $X$ and a constant $\lambda$. Any Einstein space is trivially a Ricci soliton with $X=0$. If in addition $X$ is assumed to be a non-trivial conformal vector field, then it immediately follows that the Ricci soliton is Einstein. This case is often discussed in the literature, see [12], [7]. Besides the standard sphere, compact Ricci solitons seem to be rare [5, 6]. Sophisticated analytic and topological methods have been employed for obtaining non-compact Ricci solitons [23], [3], for the approach by doubly warped products see [11].
@@
for a constant $\rho\in\mathbb{R}$ is said to evolve by the Ricci-Bourguignon flow. Here $S$ is the scalar curvature of $g=g(t).$ We will consider two particular cases. For $\rho=0$ this is the Ricci flow, for $\rho=1/n$ the right hand side equals the traceless Ricci tensor. This flow was introduced by Bourguignon [2, 3.24].
@@
Gradient Ricci-Bourguignon solitons are also named gradient $\rho$-Einstein solitons, if $\lambda$ in Equation (2) is allowed to be a function, then the metric is also called gradient Ricci-Bourguignon almost soliton. For $\lambda=0$ the gradient Ricci-Bourguignon (almost) soliton is called steady.
@@
All complete warped products with a $1$-dimensional basis satifying the equation $\textrm{Ric}+\nabla^{2}f-\frac{S}{n}g=0$ of a steady gradient Ricci-Bourguignon soliton can be classified by a family of explicitly given functions $h$ and $f$ depending on finitely many parameters.
@@ 2. Basic formulas
and a function $f(t)$ defined on the basis. Unless stated otherwise, it will be assumed that $h$ and $f$ are smooth functions and $h^{\prime}$ is not constant and non-zero. Isolated zeros of $h^{\prime}$ are allowed if the metric singularity is removable, as it is the case for polar coordinates. The gradients of $h$ and $f$ appear as $h^{\prime}\partial_{t}$ and $f^{\prime}\partial_{t}$, respectively, and the Hessian tensors are $\nabla^{2}h=h^{\prime\prime}g$ on the one hand and
@@
The Ricci tensors of $g$ and $g_{*}$ will be denoted by Ric and $\rm Ric$$_{*}$, respectively. Any equation of the type
@@
$a\textrm{Ric}(\partial_{t},\partial_{t})+bf^{\prime\prime}+c$ $=$ $0$
@@
(4) $(n-1)a\frac{h^{\prime\prime\prime}}{h^{\prime}}+bf^{\prime\prime}+c=0$
@@
Therefore one can express $f^{\prime}$ as $h^{\prime}$ times an integral of $\frac{a}{b}(n-2)\frac{h^{\prime\prime 2}-h^{\prime}h^{\prime\prime\prime}-\lambda_{*}}{h^{\prime 3}}$ and determine $f$ itself by an additional integration step, compare [19, Ex.1.5]. Finally, $c$ is obtained by the trace of $a{\rm Ric}+b\nabla^{2}f$. However, in general $c$ won’t be constant, so this method does not produce Ricci solitons but only Ricci almost solitons. Considering only the traceless part leads to the following:
@@
For any given Einstein space $M_{*}$ and any smooth function $h^{\prime}$ there is a function $f$ such that the warped product $\mathbb{R}\times_{h^{\prime}}M_{*}$ admits at least locally a solution of $(a{\rm Ric}+b\nabla^{2}f)^{\circ}=0$. For $a=b=1$ the left hand side is nothing but the traceless Bakry-Émery-Ricci tensor [22].
If $h^{\prime}$ is globally defined and everywhere strictly positive, then the warped product $\mathbb{R}\times_{h^{\prime}}M_{*}$ is complete whenever $M_{*}$ is complete. This leads to complete solutions of $({\rm Ric}+\nabla^{2}f)^{\circ}=0$ (Ricci almost solitons).
@@
    Essentially the same holds under the assumption that $a=b=1$ and $M$ is Einstein, compare [19, Thm.1.3]. Up to scaling, the list of Einstein spaces admitting a function $f$ satisfying $\nabla^{2}f=\lambda g$ for some function $\lambda$ consists of warped products $dt^{2}+(f^{\prime}(t))^{2}ds_{*}^{2}$ with a $1$-dimensional basis where $f(t)$ is a solution of the oscillator equation $f^{\prime\prime}+\sigma f=0$ where $\sigma$ is a normalized scalar curvature ($\sigma=1$ for the unit sphere) [14], [15]. Among them the unit sphere is the only compact Einstein space with $f^{\prime}(t)=\sin t$ with the two zeros added (it is not a global warped product).
@@
A solution with $a=b=1$ and a constant $c$ is normally called a gradient Ricci soliton, and the tensor ${\rm Ric}^{\infty}_{f}={\rm Ric}+\nabla^{2}f$ is called the Bakry-Émery-Ricci tensor [22], for ${\rm Ric}^{m}_{f}={\rm Ric}+\nabla^{2}f-m^{-1}df\otimes df$ see [10]. A solution with $a=b=1$ and a function $c$ is called a gradient Ricci almost soliton [19].
@@
    Any 22-dimensional gradient Ricci soliton with $b\not=0$ is a warped product outside the set of critical points which are isolated. This follows from the known solutions of the equation $\nabla^{2}f=\phi g$ with any scalar function $\phi$ [20], [14], compare [18].
@@
1. The metric $ds^{2}=dt^{2}+\tanh^{2}t\,du^{2}$ on $(t,u)\in[0,\infty)\times S^{1}$ is a well known example of a $2$-dimensional complete Ricci soliton defining a complete metric on $\mathbb{R}^{2},$ cf. Kühnel [14, Example 4(a)]. Here $c=0,\gamma=-2,h^{\prime}(t)=\tanh t,f^{\prime}=-2h^{\prime},h^{\prime\prime}(t)=(\cosh t)^{-2},K=-h^{\prime\prime\prime}/h^{\prime}=2h^{\prime\prime}=-f^{\prime\prime},\nabla^{2}f=f^{\prime\prime}g=-2h^{\prime\prime}g=K\,g.$ With $x=\sinh t\,\sin u;y=\sinh t\cos u$ this metric on $\mathbb{R}^{2}$ is of the form
@@ 3. Complete and explicit solutions
Throughout this section we consider Ricci almost solitons in a slightly generalized form as solutions of $(a\rm{Ric}+b\nabla^{2}f)^{\circ}=0$ with given constants $a$ and $b$.
@@
If $n=2$ then the metric $g_{*}$ is 11-dimensional, hence $\lambda_{*}=0,$ and the two coupled equations (4) und (6) read as:
$-ah^{\prime\prime\prime}+bh^{\prime}f^{\prime\prime}+ch^{\prime}$ $=$ $0$
$ah^{\prime}h^{\prime\prime\prime}-bf^{\prime}h^{\prime}h^{\prime\prime}-ch^{\prime 2}$ $=$ $0$
@@
(11) $-ah^{\prime\prime\prime}+b\gamma h^{\prime}h^{\prime\prime}+ch^{\prime}$ $=$ $0$
@@
Proof. The sum of both equations (the second divided by $h^{\prime}$) leads to $b(h^{\prime}f^{\prime\prime}-h^{\prime\prime}f^{\prime})=0$, so we have either b=0 or $\log(f^{\prime})^{\prime}=\log(h^{\prime})^{\prime}$ which is equivalent to $f^{\prime}=\gamma h^{\prime}$ for a constant $\gamma.$ Then both equations are converted into (11). For $a=1,b=0$ see Prop. 2.2 (1). $\Box$
There is a method for solving an ODE of type $y^{\prime\prime}=F(y,y^{\prime})$ [21, §11]. We concentrate on the case $a=b=1$.
@@
For $c=0$ Equation (11) is of first order by $(h^{\prime\prime}-\frac{\gamma}{2}h^{\prime 2})^{\prime}=0$. There are various odd functions $h^{\prime}$ solving this equation. Typical solutions with initial condition $h^{\prime}(0)=0$ are $h^{\prime}(t)=\tan t$ for $\gamma=2$ and $h^{\prime}(t)=\tanh t=\frac{e^{2t}-1}{e^{2t}+1}$ for $\gamma=-2$ (the complete cigar metric.) Other typical solutions are $h^{\prime}(t)=\frac{1}{t}$ for $\gamma=-2$ and $h^{\prime}(t)=\coth t=\frac{e^{2t}+1}{e^{2t}-1}$ for $\gamma=-2$. There are no even functions solving this equation.
@@
Let $h(t)=\arctan(t)$, then for any $\lambda_{*}$ there is a global function $f(t)$ such that $\mathbb{R}\times_{h^{\prime}}M_{*}$ is a complete Ricci almost soliton whenever $M_{*}$ is complete with Einstein constant $\lambda_{*}$.
@@
$(t^{2}+1)f^{\prime}=\frac{a\lambda_{*}}{b}\int(t^{2}+1)^{3}dt+\frac{a(n-2)}{b}\int\frac{2(t^{2}-1)}{t^{2}+1}dt$
$\frac{a\lambda_{*}}{b}\Big(\frac{t^{7}}{7}+\frac{3t^{5}}{5}+t^{3}+t\Big)+\frac{a(n-2)}{b}\Big(2t-4\arctan(t)+C\Big).$
@@
If $g_{*}$ is the standard metric on the sphere, then the space can be viewed as a rotationally symmetric deformed infinite cylinder with a “cusp” at $\pm\infty$. This satisfies the equation
@@
Let $h(t)=\frac{1}{3}t^{3}+t$, then for any $\lambda_{*}$ there is a global function $f(t)$ such that $\mathbb{R}\times_{h^{\prime}}M_{*}$ is a complete Ricci almost soliton whenever $M_{*}$ is complete with Einstein constant $\lambda_{*}$.
@@
Let $h(t)=(t^{2}+2)e^{t}$, then for any $\lambda_{*}$ there is a global function $f(t)$ such that $\mathbb{R}\times_{h^{\prime}}M_{*}$ is a complete Ricci almost soliton whenever $M_{*}$ is complete with Einstein constant $\lambda_{*}$.
@@
$2a(n-2)(t^{2}+12t+8)e^{2t}+b(t^{2}+2t+2)^{3}e^{3t}\Big(\frac{f^{\prime}}{(t^{2}+2t+2)e^{t}}\Big)^{\prime}=a\lambda_{*}.$
@@ 4. Ricci-Bourguignon solitons
with a constant $\lambda$ where $S$ denotes the scalar curvature [7]. The case $\lambda=0$ is then referred to as a steady gradient Ricci-Bourguignon soliton.
@@
2. If the gradient of $f$ is assumed to be a conformal vector field, then Equation (12) implies that $M$ is Einstein. Consequently, even if $\lambda$ is admitted to be a non-constant function, the only complete cases are listed in [14, Thm.27], implicitly already in [20, Sect.3]. In particular, the only compact case is the standard sphere. Implicitly this result can already be found in [20, Thm.1] since an Einstein space conformal to the standard sphere is isometric to the standard sphere. It is restated in [7, Thm.1.5]. For the last statement a constant scalar curvature is sufficient [14, Thm.24].
3. Complete gradient Ricci-Bourguignon almost solitons can be constructed by the same principle as in Section 3, compare [7, Ex.1.7].
@@
A warped product $M=\mathbb{R}\times_{h^{\prime}}M_{*}$ of dimension $n\geq 4$ with a potential function $f(t)$ is a steady gradient Ricci-Bourguignon soliton if and only if $f$ is harmonic and Equation (8) is satisfied with $a=b=1$.
@@
$ds^{2}=dt^{2}+(h^{\prime}(t))^{2}ds_{*}^{2}$ is a steady gradient Ricci-Bourguignon soliton if and only if $h$ and $f$ satisfy the following equations with the constant $\lambda_{*}$ as the normalized scalar curvature of $g_{*}$:
@@
(14) $\Delta f=f^{\prime\prime}+(n-1)\frac{f^{\prime}h^{\prime\prime}}{h^{\prime}}$ $=$ $0.$
@@
$(n-2)\big[h^{\prime\prime 2}-h^{\prime}h^{\prime\prime\prime}\big]+h^{\prime}\big[f^{\prime\prime}h^{\prime}-f^{\prime}h^{\prime\prime}\big]$ $=$ $0$
$f^{\prime\prime}+(n-1)\frac{f^{\prime}h^{\prime\prime}}{h^{\prime}}$ $=$ $0.$